Why is the constant K in coulumbs law is eqaul to 1/4pie epsilon not......

and second que is that
why only K is used as constant of proportionality in coulumbs Law..??
1
Expert's answer
2019-09-04T09:22:38-0400

1.This is because of gauss law which state that

close integral of Electrical field "E" over area is equal to "\frac{q}{\epsilon}"

So,

Considering surface to be spherical

"\overset{s_1}{\int} E.ds=\frac{q}{\epsilon}=E\overset{s_1}{\int} ds"

"s_1=4\pi r^2"

"E.4\pi r^2=\frac{q}{\epsilon}"

"E=\frac{1}{4\pi \epsilon}\frac{q}{r^2}"

"F=q_{test}E=\frac{1}{4\pi \epsilon}\frac{q_{test}q}{r^2}"

"k=\frac{1}{4\pi \epsilon}"

So, value of "k" "=\frac{1}{4\pi \epsilon}"


2.

Let 2 charges "q_1,q_2" are seperate by "r" distance

Accrding to coulomb law

"F" is directly proportional to "q_1q_2"

"F" is directly proportional to "\frac{1}{r^2}"

combining both

"F" is directly proportional to "\frac{q_1q_2}{r^2}"

removing proportionality sign by putting a constant "k"

"F=\frac{kq_1q_2}{r^2}"

So. only "k" is used as constant of proportionality in coulomb law